# Delving Deeper Into Astromorphic Transformers

### v2

@@ Abstract
Preliminary attempts at incorporating the critical role of astrocytes - cells that constitute more than 50% of human brain cells - in brain-inspired neuromorphic computing remain in infancy. This paper seeks to delve deeper into various key aspects of neuron-synapse-astrocyte interactions to mimic self-attention mechanisms in Transformers. The cross-layer perspective explored in this work involves bio-plausible modeling of Hebbian and pre-synaptic plasticities in neuron-astrocyte networks, incorporating effects of non-linearities and feedback along with algorithmic formulations to map the neuron-astrocyte computations to self-attention mechanism and evaluating the impact of incorporating bio-realistic effects from the machine learning application side. Our analysis on sentiment and image classification tasks on the IMDB and CIFAR10 datasets underscores the importance of constructing Astromorphic Transformers from both accuracy and learning speed improvement perspectives.
@@ I Introduction
Astrocytes, a type of glial cell, play a critical role in brain function encompassing various processes such as homeostasis, metabolism, and synaptic regulation [1]. Astrocytes detect and regulate synaptic activity in the tripartite synapse through interactions with pre and postsynaptic neurons. Investigating their impact on neural computation is currently an active research field in neuroscience and underscores the critical need to move beyond the neuro-synaptic perspective of current Artificial Intelligence (AI) systems. Recent experimental findings on neuron-astrocyte interactions and modulation have led to significant progress in computational neuroscience, enabling the development of models that incorporate neuron-astrocyte interactions within neural networks [2, 3]. Astrocytes have been found to modulate bursting in neural circuitry through the release of gliotransmitters, which have an impact on neuronal excitability and synaptic plasticity [4, 5]. Astrocytes possess the ability to encode information through calcium signaling and regulate information processing, thereby actively engaging in neural computation at the tripartite synapse level. Additionally, astrocytes possess inherent capacity as memory components [6, 7] and plasticity regulators that are capable of facilitating local sequential learning [8, 9].
@@
Preliminary attempts at incorporating the critical role of astrocytes in brain-inspired neuromorphic computing has started in earnest. Neuromorphic systems [10] are bio-plausible AI models that attempt to emulate neuronal and synaptic operations with a higher degree of bio-fidelity to tap into the huge energy efficiency of the brain. Such systems are temporal and event-driven where computation occurs through transmission of sparse spiking events between neurons - thereby offering the potential of orders of magnitude power efficiency when implemented on neuromorphic platforms [11]. However, astromorphic computing encapsulating features of astrocytic regulation in neuromorphic systems still remain in infancy with works primarily targeting integration of astrocyte functionality with Spike-Timing-Dependent-Plasticity (STDP) mechanisms [12, 13, 14, 15] - an unsupervised learning framework that remains limited to shallow networks and simple tasks [16]. Other attempts also suffer from the pressing issue of scalability to complex tasks [17].
@@
This work is based on a recent study [18] that has shown promise at emulating the self-attention mechanism in transformers [19] using a simple two layered network involving neuron-astrocyte interactions. The ability to design Astromorphic Transformers is therefore highly relevant in the current era of Large Language Models (LLMs), such as BERT [20] and GPT [21] - paving the pathway for designing explainable transformers from a neuroscience perspective as well as addressing scaling challenges for utilizing the computational role of astrocytes in neuromorphic computing. However, as we illustrate in this work, prior work has neglected various key aspects of bi-directional astrocytic signalling, feedback and non-linear temporal behavior - all of which are key to address the fundamental mapping of self-attention mechanism to astrocyte-neuron interactions along with providing insights on whether such intrinsic non-linear dynamics and plasticity modulation enable better learning capabilities from the application side. The distinguishing feature of this work therefore lies in the neuroscience-algorithm-application co-design perspective seeking to answer the following questions: (i) Theoretical neuroscience: What aspects of astrocyte functionality should be included in neural architectures to better explain self-attention mechanism in transformers? (ii) Algorithm: Can we map key algorithmic modules like relative positional information encoding of tokens to astrocytic interactions by drawing qualitative inspiration from theoretical neuroscience? (iii) Application: What is the impact of incorporating such dynamic temporal bio-plausible transformers from the machine learning application perspective?
@@ II Related Works and Main Contributions
Understanding the role of astrocytes in neuromorphic computing, especially from the algorithm perspective, is relatively nascent. While hardware implementation of astrocyte functionality in CMOS [22, 23] as well as post-CMOS technologies [24] have been explored, algorithm level works exploiting the role of astrocytes and leveraging their computational role has largely remain limited to self-repair functionality of faulty neural network hardware [25, 12, 26, 14]. This is mainly inspired by the fact that astrocytes, through their synaptic plasticity regulation and feedback mechanisms, enable neurons to fire at a baseline ideal frequency. This property can be utilized to mitigate neuromorphic hardware non-idealities with stuck-at-faults and drift effects [14]. However, such approaches are still limited to single layered unsupervised networks where astrocyte functionality is integrated with STDP learning mechanisms.
Recent studies have also examined the influence of astrocytes on learning [9], working memory [6], energy minimization in neural networks [27], robotic locomotion control [28], structure learning [17], among others. Building upon recent works [18] at emulating self-attention module in transformers through astrocyte-neuron interactions, we develop a computational framework that encapsulates key aspects of temporal non-linearities and feedback enabled by astrocytes and underscore the importance of incorporating such bio-plausible effects through a comprehensive evaluation of Astromorphic Transformers on a benchmark evaluation set. Our work expands the computational capabilities of astrocyte-neuron network architectures equipped with plasticity, for the first time, to a broad category of sequential tasks.
@@ III-A Transformers
Vanilla transformers [19] for classification tasks usually consist of an encoder block processing the provided key$(K)$, query$(Q)$ and value$(V)$ generated from the input token embeddings $X$ (words in language tasks, pixels in vision tasks, etc.). Appendix A in the supplementary material illustrates the conventional configuration of a transformer, featuring a block referred to as “Multi-Head Attention” that produces the contextual association among the provided inputs. The Multi-Head Attention block in transformers uses self-attention mechanisms for multiple input heads, allowing the model to allocate varying importance to individual tokens, focusing on salient features for precise output in specific contexts.
The transformer implements auto-regressive attention by computing the dot product of key and query, and then multiplying the resulting scalar with the corresponding value derived from the input token. Implementing the softmax function (a type of distribution function) atop the dot product of key and query (to enforce attention over the tokens) increases the order of complexity, a predicament that can be alleviated by the use of linearized attention [29]. Linearized attention exploits the distributive property of matrix multiplication resulting in lower complexity for self-attention computation. More details can be found in Appendix B of the supplementary material. From the linear attention-based transformer (linear transformer) perspective [30], the transformer equations can be defined in the following form:
@@
$Z=\mathrm{Softmax}(\mathrm{Linear}(\mathrm{LayerNorm}(\mathrm{FFN}(Y)+Y)))$ (3)
From Equations 2 and 3, we can see that the output of self-attention is $Y$, which undergoes layer normalization after being augmented with residual connections from the input layer. The transformer determines the final output probabilities, denoted by $Z$. There is a $\phi(.)$ feature map that corresponds to a positive similarity function. When two such functions are multiplied, the normalized dot product maps to a distribution function [31] (i.e. softmax). According to [30], $\phi(.)$ can be defined by Equation 4. Here, the activation function $elu(x)$, which stands for exponential linear unit, is utilized for efficient and precise learning [32].
$\phi(x)=\mathrm{elu}(x)+1;$ (4)
In the next section, we discuss the tripartite synapse model and aspects which will be necessary in constructing astromorphic transformers.
@@ III-B1 Astrocyte dynamics
Figure 1: A model of synaptic communication in the brain. The tripartite synapse consists of presynaptic neurons, postsynaptic neurons and astrocytes. Astrocytes detect neuronal activity and respond bidirectionally by emitting gliotransmitters, thereby modulating the intensity and duration of synaptic communication.
According to the biophysical models, it is observed that an action potential (AP) is generated by the presynaptic axon through the release of neurotransmitters. Neurotransmitters bind to postsynaptic dendrites, causing depolarization. This leads to the release of 2-Arachidonylglycerol (2-AG) from postsynaptic neurons to type 1 Cannabinoid Receptors (CB1Rs) on the presynaptic terminal. This process is called Depolarization-induced Suppression of Excitation ($DSE$) and depresses the continuous release of neurotransmitters by inhibiting presynaptic neurons [38]. Neurotransmitter release also involves binding 2-AG to CB1Rs on an astrocyte, enveloping the synapse and increasing Inositol 1,4,5-trisphosphate ($IP_{3}$) levels. $IP_{3}$ facilitates the release of calcium ions into the cytoplasm, thereby, increasing calcium concentration. Three currents, $J_{channel}$, $J_{leak}$, and $J_{pump}$, are introduced to model the rate of change of calcium concentration inside the astrocyte [33]. When calcium concentration crosses a threshold, astrocytes discharge glutamate, which binds to presynaptic group I metabotropic Glutamate Receptors (mGluRs), creating a process called eSP (Indirect Feedback) from the astrocyte to the presynaptic neuron [39]. The comprehensive description of the computational model (Equation 5 introduces the basic equation of the model) can be found in the works of [40, 37, 41].
@@ III-B2 Neuron dynamics
In this work, we consider the passive Leaky Integrate and Fire (LIF) dynamics to describe the neuron firing process, which is described as follows:
@@
where, $\tau_{m}$ represents the membrane time constant, $v$ represents the membrane potential, and $R_{m}$ represents the membrane resistance. $I_{total-syn}^{i}$ denotes the total synaptic current at the $i^{th}$ synapse. $M$ is the number of synapses connecting to the corresponding neuron. Total synaptic current at the $i^{th}$ synapse can be defined as the summation of two currents namely, Synaptic Current, $I_{syn}^{i}$ and Slow Inward Current, $SIC(t)$ in Equation 7:
@@
$I_{syn}^{i}$ is the intrasynaptic current which varies linearly with the amount of neurotransmitters present in that synapse at any time, $t$ ($I_{syn}^{i}=A_{se}y_{i}(t)$ where, $A_{se}$ and $y_{i}(t)$ represent the absolute synaptic efficacy and the amount of neurotransmitter discharged by the synapse at a given time $t$, respectively). Thus, $I_{syn}$ depends on the direct feedback (corresponds to $DSE$) as direct feedback is modulated by the amount of neurotransmitters between the pre and postsynaptic neuron. On the other hand, the extrasynaptic Slow Inward Current $(SIC(t))$ is generated when astrocyte-released glutamate interacts with extrasynaptic NMDARs (located in postsynaptic neuron) [42]. To elaborate, the activation of astrocytes leads to the release of glutamate from these cells. Upon its release, glutamate interacts with NMDARs located on the postsynaptic neuron, resulting in the emergence of a distinct form of current known as Slow Inward Current $(SIC(t))$ [37]. The details of the mathematical modeling of $SIC(t)$ can be found in [37].
Next, we develop a network architectural framework to mimic the self-attention mechanism in transformers using the astrocyte and neuron dynamics described herein and evaluate the impact of inherent non-linearities and feedback.
Figure 2: The neural network architecture showing the three layers. As tokens are presented to the network as a $d$ dimensional vector, there are $d$ neurons in both the input and output layers. Hidden layer has $m$ neurons. Solid lines indicate active operations, therefore $H_{neuron}$ and $H_{astro}$ are learnt during write mode but utilized during read mode.
@@ III-C Neural Network Emulating Self-Attention
Self-attention is a fundamental component of all transformer models. Based on prior works of mimicking self-attention by astrocytes [18], we adopt a network topology as shown in Fig. 2. The neural network architecture comprises three distinct layers, namely the Input layer, Hidden layer, and Output layer. There exists two distinct operational modes for the aforementioned neural network: (i) write mode and (ii) read mode. Upon presentation of the tokens ($x_{t}\in R^{1\times d}$: which may include word embeddings and image embeddings) to the input layer, the write mode is initially employed, followed by the read mode to generate the outputs. In write mode, we encode information into the network by means of keys and values, whereas, in read mode, we retrieve the information by providing queries.
During the write mode, the input layer sequentially transmits the keys ($k_{t}=x_{t}W_{K};\;W_{K}\in R^{d\times m}$) for each token to the hidden layer and the values ($v_{t}=x_{t}W_{V};\;W_{V}\in R^{d\times d}$) to the output layer. The hidden layer neurons introduce a nonlinearity within the system by activating a feature map, denoted as $\phi(.)$, on the transmitted keys. Use of $\phi(.)$ reduces the computational complexity and memory usage inspired from linearized attention mechanism. More information on the reduction of the order of complexity is discussed in Appendix B (with Fig. S2 illustrating the idea) in the supplementary material. A presynaptic neuron in the hidden layer is intricately connected with an astrocyte and an output layer neuron, forming a tripartite synapse. Similar to the mechanism by which the presynaptic neuron stimulates the astrocyte through the release of neurotransmitters, in this context, the astrocyte is being activated through the transmission of keys from neurons in the hidden layer. The write mode can be mathematically expressed as in Equation 8:
$f_{t}=x_{t}\qquad\qquad\qquad\;\;\;$ (8)
@@
$l_{t}=f_{t}W_{V}=v_{t}\qquad\quad$
In order to establish the read mode equations, it is necessary to first formulate the neural plasticities that occur among the components of the tripartite synapse. Moreover, our work aims at exploring strong bio-plausible correlations at modelling such plasticities, which ultimately influence our neuroscience-algorithm-application co-design analysis. In a subsequent section, we will examine the read mode which is responsible for retrieving the encoded information.

### v3

@@ Abstract
Preliminary attempts at incorporating the critical role of astrocytes—cells that constitute more than 50% of human brain cells—in brain-inspired neuromorphic computing remain in infancy. This paper seeks to delve deeper into various key aspects of neuron-synapse-astrocyte interactions to mimic self-attention mechanisms in Transformers. The cross-layer perspective explored in this work involves bioplausible modeling of Hebbian and presynaptic plasticities in neuron-astrocyte networks, incorporating effects of non-linearities and feedback along with algorithmic formulations to map the neuron-astrocyte computations to self-attention mechanism and evaluating the impact of incorporating bio-realistic effects from the machine learning application side. Our analysis on sentiment and image classification tasks (IMDB and CIFAR10 datasets) highlights the advantages of Astromorphic Transformers, offering improved accuracy and learning speed. Furthermore, the model demonstrates strong natural language generation capabilities on the WikiText-2 dataset, achieving better perplexity compared to conventional models, thus showcasing enhanced generalization and stability across diverse machine learning tasks.
@@ I Introduction
Astrocytes, a type of glial cell, play a critical role in brain function, encompassing various processes such as homeostasis, metabolism, and synaptic regulation [1]. Astrocytes detect and regulate synaptic activity in the tripartite synapse through interactions with pre- and postsynaptic neurons. Investigating their impact on neural computation is currently an active research field in neuroscience and underscores the critical need to move beyond the neuro-synaptic perspective of current Artificial Intelligence (AI) systems. Recent experimental findings on neuron-astrocyte interactions and modulation have led to significant progress in computational neuroscience, enabling the development of models that incorporate neuron-astrocyte interactions within neural networks [2, 3]. Astrocytes have been found to modulate bursting in neural circuitry through the release of gliotransmitters, which have an impact on neuronal excitability and synaptic plasticity [4, 5]. Astrocytes possess the ability to encode information through calcium signaling and regulate information processing, thereby actively engaging in neural computation at the tripartite synapse level. Additionally, astrocytes possess inherent capacity as memory components [6, 7] and plasticity regulators that are capable of facilitating local sequential learning [8, 9].
@@
Preliminary attempts at incorporating the critical role of astrocytes in brain-inspired neuromorphic computing have started in earnest. Neuromorphic systems [10] are bioplausible AI models that attempt to emulate neuronal and synaptic operations with a higher degree of bio-fidelity to tap into the huge energy efficiency of the brain. Such systems are temporal and event-driven, where computation occurs through transmission of sparse spiking events between neurons, thereby offering the potential of orders of magnitude power efficiency when implemented on neuromorphic platforms [11]. However, astromorphic computing that encapsulates the features of astrocytic regulation in neuromorphic systems still remains in infancy, with works focusing primarily on the integration of astrocyte functionality with Spike-Timing-Dependent-Plasticity (STDP) mechanisms [12, 13, 14, 15]—an unsupervised learning framework that remains limited to shallow networks and simple tasks [16]. Other attempts also suffer from the pressing issue of scalability to complex tasks [17].
@@
This work is based on a recent study [18] that has shown promise at emulating the self-attention mechanism in transformers [19] using a simple two-layered network involving neuron-astrocyte interactions. The ability to design Astromorphic Transformers is therefore highly relevant in the current era of Large Language Models (LLMs), such as BERT [20] and GPT [21], paving the pathway for designing explainable transformers from a neuroscience perspective as well as addressing scaling challenges for utilizing the computational role of astrocytes in neuromorphic computing. However, as we illustrate in this work, prior work has neglected various key aspects of bi-directional astrocytic signalling, feedback, and non-linear temporal behavior—all of which are key to addressing the fundamental mapping of self-attention mechanism to astrocyte-neuron interactions along with providing insights on whether such intrinsic non-linear dynamics and plasticity modulation enable better learning capabilities from the application side. The distinguishing feature of this work, therefore, lies in the neuroscience-algorithm-application co-design perspective seeking to answer the following questions: (i) Theoretical neuroscience: What aspects of astrocyte functionality should be included in neural architectures to better explain the self-attention mechanism in transformers? (ii) Algorithm: Can we map key algorithmic modules, like the relative positional information encoding of tokens, to astrocytic interactions by drawing qualitative inspiration from theoretical neuroscience? (iii) Application: What is the impact of incorporating such dynamic temporal bioplausible transformers from the machine learning application perspective?
@@ II Related Works and Main Contributions
Spiking neural networks (SNNs), inspired by biological neurons, have garnered attention for their ability to process temporal data efficiently through sparse, event-driven computation and communication [22, 23]. Recent works have explored enhancing the computational efficiency of SNNs, including incorporation of advanced learning techniques [25], improving network robustness through hybrid models [26], and addressing energy efficiency challenges [27, 28]. While these advances contribute to energy-efficient and robust learning systems, they remain focused on spike-based computation and do not fully leverage neuron-astrocyte interactions, which is the core focus of our work.
Understanding the role of astrocytes in neuromorphic computing, especially from the algorithm perspective, is relatively nascent. While hardware implementation of astrocyte functionality in CMOS [29, 30] as well as post-CMOS technologies [31] have been explored, algorithm-level works exploiting the role of astrocytes and leveraging their computational role have largely remained limited to self-repair functionality of faulty neural network hardware [32, 12, 33, 14]. These works are mainly inspired by the fact that astrocytes, through their synaptic plasticity regulation and feedback mechanisms, enable neurons to fire at a baseline ideal frequency. This property can be utilized to mitigate neuromorphic hardware non-idealities with stuck-at-faults and drift effects [14]. However, such approaches are still limited to single-layered unsupervised networks where astrocyte functionality is integrated with STDP learning mechanisms. Recent studies have also examined the influence of astrocytes on learning [9], working memory [6], energy minimization in neural networks [34], robotic locomotion control [35], structure learning [17], among others.
Building upon recent works [18] at emulating self-attention module in transformers through astrocyte-neuron interactions, we develop a computational framework that encapsulates key aspects of temporal non-linearities and feedback enabled by astrocytes and underscore the importance of incorporating such bioplausible effects through a comprehensive evaluation of Astromorphic Transformers on a benchmark evaluation set. Our work expands the computational capabilities of astrocyte-neuron network architectures equipped with plasticity, for the first time, to a broad category of sequential tasks.
@@ III-A Transformers
Vanilla transformers [19] for classification tasks usually consist of an encoder block processing the provided key$(K)$, query$(Q)$, and value$(V)$ generated from the input token embeddings $X$ (words in language tasks, pixels in vision tasks, etc.). Appendix A in the supplementary material illustrates the conventional configuration of a transformer, featuring a block referred to as “Multi-Head Attention” that produces the contextual association among the provided inputs. The multi-head attention block in transformers uses a self-attention mechanism for multiple input heads, allowing the model to allocate varying importance to individual tokens, focusing on salient features for precise output in specific contexts.
The transformer implements auto-regressive attention by computing the dot product of key and query and then multiplying the resulting scalar with the corresponding value derived from the input token. Implementing the softmax function (a type of distribution function) atop the dot product of key and query (to enforce attention over the tokens) increases the order of complexity, a predicament that can be alleviated by the use of linearized attention [36]. Linearized attention exploits the distributive property of matrix multiplication, resulting in lower complexity for self-attention computation. More details can be found in Appendix B of the supplementary material. From the linear attention-based transformer (linear transformer) perspective [37], the transformer equations can be defined in the following form:
@@
$\mathrm{logits}=\mathrm{Softmax}(\mathrm{Linear}(\mathrm{LayerNorm}(\mathrm{FFN}(Y)+Y)))$ (3)
From Equations 2 and 3, we can see that the output of self-attention is $Y$, which undergoes layer normalization after being augmented with residual connections from the input layer. The transformer determines the final output probabilities, denoted by logits. $\mathrm{FFN}$ stands for a Feed Forward Network and there is a $\phi(.)$ feature map that corresponds to a positive similarity function. When two such functions are multiplied, the normalized dot product maps to a distribution function [38] (i.e., softmax). According to [37], $\phi(.)$ can be defined by Equation 4. Here, the activation function includes $elu(x)$, which stands for exponential linear unit and is utilized for efficient and precise learning [39].
$\phi(x)=\mathrm{elu}(x)+1$ (4)
In the next section, we discuss the tripartite synapse model and aspects that will be necessary in constructing Astromorphic Transformers.
@@ III-B1 Astrocyte dynamics
Figure 1: A model of synaptic communication in the brain. The tripartite synapse consists of presynaptic neurons, postsynaptic neurons, and astrocytes. Astrocytes detect neuronal activity and respond bidirectionally by emitting gliotransmitters, thereby modulating the intensity and duration of synaptic communication.
According to the biophysical models, it is observed that an action potential (AP) is generated by the presynaptic axon through the release of neurotransmitters. Neurotransmitters bind to postsynaptic dendrites, causing depolarization. This leads to the release of 2-Arachidonoylglycerol (2-AG) from postsynaptic neurons to type 1 Cannabinoid Receptors (CB1Rs) on the presynaptic terminal. This process is called Depolarization-induced Suppression of Excitation ($DSE$) and depresses the continuous release of neurotransmitters by inhibiting presynaptic neurons [45]. Neurotransmitter release also involves binding 2-AG to CB1Rs on an astrocyte, enveloping the synapse, and increasing Inositol 1,4,5-trisphosphate ($IP_{3}$) levels. $IP_{3}$ facilitates the release of calcium ions into the cytoplasm, thereby increasing calcium concentration. Three currents, $J_{channel}$, $J_{leak}$, and $J_{pump}$, are introduced to model the rate of change of calcium concentration inside the astrocyte [40]. When calcium concentration crosses a threshold, astrocytes discharge glutamate, which binds to presynaptic group I metabotropic Glutamate Receptors (mGluRs), creating a process called eSP (Indirect Feedback) from the astrocyte to the presynaptic neuron [46]. The comprehensive description of the computational model (Equation 5 introduces the basic equation of the model) can be found in the works of [47, 44, 48], which are outlined in Appendix D in the supplementary material.
@@ III-B2 Neuron dynamics
In this work, we consider the passive Leaky Integrate and Fire (LIF) dynamics to describe the neuron firing process, which is described in Equation 6:
@@
where $\tau_{m}$ represents the membrane time constant, $v$ represents the membrane potential, and $R_{m}$ represents the membrane resistance. $I_{total-syn}^{i}$ denotes the total synaptic current at the $i^{th}$ synapse. $M$ is the number of synapses connecting to the corresponding neuron. Total synaptic current at the $i^{th}$ synapse can be defined as the summation of two currents, namely, Synaptic Current, $I_{syn}^{i}$ and Slow Inward Current, $SIC(t)$ in Equation 7:
@@
$I_{syn}^{i}$ is the intrasynaptic current, which varies linearly with the amount of neurotransmitters present in that synapse at any time, $t$ ($I_{syn}^{i}=A_{se}y_{i}(t)$, where $A_{se}$ and $y_{i}(t)$ represent the absolute synaptic efficacy and the amount of neurotransmitter discharged by the synapse at a given time $t$, respectively). Thus, $I_{syn}$ depends on the direct feedback (corresponds to $DSE$) as direct feedback is modulated by the amount of neurotransmitters between the pre- and postsynaptic neuron. On the other hand, the extrasynaptic Slow Inward Current $(SIC(t))$ is generated when astrocyte-released glutamate interacts with extrasynaptic NMDARs (located in postsynaptic neurons) [49]. To elaborate, the activation of astrocytes leads to the release of glutamate from these cells. Upon its release, glutamate interacts with NMDARs located on the postsynaptic neuron, resulting in the emergence of a distinct form of current known as Slow Inward Current $(SIC(t))$ [44]. The details of the mathematical modeling of $SIC(t)$ can be found in [44]. Next, we develop a network architectural framework to mimic the self-attention mechanism in transformers using the astrocyte and neuron dynamics described herein and evaluate the impact of inherent non-linearities and feedback.
Figure 2: The neural network architecture showing the three layers. As tokens are presented to the network as a $d$ dimensional vector, there are $d$ neurons in both the input and output layers. The hidden layer has $m$ neurons. Solid lines indicate active operations; therefore, $H_{neuron}$ and $H_{astro}$ are learned during write mode but utilized during read mode.
@@ III-C Neural Network Emulating Self-Attention
Self-attention is a fundamental component of all transformer models. Based on prior works of mimicking self-attention by astrocytes [18], we adopt a network topology as shown in Fig. 2. The neural network architecture comprises three distinct layers, namely input layer, hidden layer, and output layer. There exist two distinct operational modes for the aforementioned neural network: (i) write mode and (ii) read mode. Upon presentation of the tokens ($x_{t}\in R^{1\times d}$: which may include word embeddings and image embeddings) to the input layer, the write mode is initially employed, followed by the read mode to generate the outputs. In write mode, we encode information into the network by means of keys and values, whereas, in read mode, we retrieve the information by providing queries.
During the write mode, the input layer sequentially transmits the keys ($k_{t}=x_{t}W_{K};\;W_{K}\in R^{d\times m}$) for each token to the hidden layer and the values ($v_{t}=x_{t}W_{V};\;W_{V}\in R^{d\times d}$) to the output layer. The hidden layer neurons introduce a non-linearity within the system by activating a feature map, denoted as $\phi(.)$, on the transmitted keys. The use of $\phi(.)$ reduces the computational complexity and memory usage inspired from the linearized attention mechanism. More information on the reduction of the order of complexity is discussed in Appendix B (with Fig. S2 illustrating the idea) in the supplementary material. A presynaptic neuron in the hidden layer ($h_{t}\in R^{1\times m}$) is intricately connected with an astrocyte and an output layer ($l_{t}\in R^{1\times d}$) neuron, forming a tripartite synapse. Similar to the mechanism by which the presynaptic neuron stimulates the astrocyte through the release of neurotransmitters, in this context, the astrocyte is being activated through the transmission of keys ($k_{t}\in R^{1\times m}$) from neurons in the hidden layer. The write mode can be mathematically expressed as in Equation 8:
$f_{t}=x_{t}$ (8)
@@
$l_{t}=f_{t}W_{V}=v_{t}$
In order to establish the read mode equations, it is necessary to first formulate the neural plasticities that occur among the components of the tripartite synapse. Moreover, our work aims at exploring strong bioplausible correlations in modeling such plasticities, which ultimately influence our neuroscience-algorithm-application co-design analysis. In a subsequent section, we will examine the read mode, which is responsible for retrieving the encoded information.